# Many-body quantum dimerization in 2D atomic arrays

### v1

@@ II.3 State characterization
Figure 4: (a) Absolute value of the spin-spin correlation function for two lattices of sizes $N=2\times 10$ and $N=5\times 4$. In the first geometry we fixed: $\mathbf{i}=(1,1)$ and $\bm{\ell}=(1,\ell)$ while in the second $\mathbf{i}=(2,1)$ and $\bm{\ell}=(2,\ell)$. (b) Bipartite entanglement entropy in a ladder as a function of system size for two different partitions as indicated in the cartoon sketches. (c) Average nearest-neighbor concurrence as a function of inter-atomic distance for a $N=4\times 4$ lattice. In all plots, we considered the two models given in Eqs. (3) and (4) fixing the inter-atomic distance to $k_{0}d=0.1\pi$ and $k_{0}d=0.42\pi$, respectively. The results obtained from the RVB ansatz are also reported for direct comparison.
@@
    where $\lambda_{i}$ are the eigenvalues in decreasing order of the operator $R=\sqrt{\sqrt{\varrho}\tilde{\varrho}\sqrt{\varrho}}$ with $\tilde{\varrho}=(\sigma_{y}\otimes\sigma_{y})\varrho^{*}(\sigma_{y}\otimes\sigma_{y})$ and $\varrho$ is the corresponding density matrix. This measurement ranges from 00 for non-entangled states to $1$ for maximally entangled configurations. Since an RVB liquid ansatz consists of a covering of nearest-neighbor dimers, we expect for this state the two-qubit concurrence to be finite only between nearest-neighbor sites. To quantify this, we define the average nearest-neighbour concurrence
@@
    where $N_{\langle ij\rangle}$ is the number of nearest neighbour atoms and $C(\varrho_{\langle ij\rangle})$ is the two-qubit concurrence computed for the density operator reduced to a given nearest neighbour pair $\langle ij\rangle$. The dependence of this quantity from the inter-atomic distance is shown in Fig. 4(c). The figure shows that there exists broad ranges of interatomic distances where the average nearest-neighbor concurrence remains close to the value expected from the RVB ansatz (dashed line), indicating the formation of a dimer-like entanglement structure within the system.
@@ III Steady state preparation
Figure 5: Scaling with system size of the decay rate of the least radiant state, $\gamma_{s=1}$, for model (3) (first panel) and model (4) (second panel). In the first case, the interatomic distance is fixed at $k_{0}d=0.1\pi$, while in the second case, it is set to $k_{0}d=0.42\pi$. In both cases, we compare two different lattice geometries and include the decay rate scaling associated with $N/2$ dimers.
@@
As previously discussed, the least radiant state at half-filling is an eigenstate of the non-Hermitian Hamiltonian given in Eq. (2) and thus it has a finite decay rate. The scaling of this decay rate with the system size for the two cases in Eqs. (3)-(4) is shown in Fig. 5 for two different geometrical configurations. In all cases, we observe that the decay rate is upper-bounded by the single dimer decay rate times the number of excited emitters, $\gamma_{D}N/2$, where $\gamma_{D}$ is the decay rate of a single dimer, and remains relatively close to this bound. This implies that as the system size increases, even though this state remains the least radiant within the half-filling subspace, its decay rate can surpass that of an independent single emitter. This finding presents a significant challenge for preparing this state via a relaxation process. An initial half-filled configuration will rapidly radiate into excitation subspaces with lower filling, and the desired state will only be prepared with high probability under the condition that $N/2$ excitations remain throughout the system’s evolution, an event that becomes exponentially unlikely in time.
@@
Another strategy for preparing the dimerized state involves employing a steady-state protocol that targets the state by matching the energy and symmetry of the many-body wavefunction with the frequency and symmetry pattern of a driving field [92, 93, 94]. More specifically, we assume that each atom in the array can be individually driven by a coherent field. This is incorporated by adding the following driving Hamiltonian (in the rotating wave approximation) to the master equation given in Eq. (1):
@@
where $\Omega_{i}$ is the driving strength on the $i$-th atom and $\delta_{i}$ is the detuning between the laser and the $i$-th atomic frequency. The pattern of the driving strength is chosen to have staggered (opposite) phases over the lattice to match the symmetry of the target many-body wave-function. For a finite-sized array, dipole-dipole interactions give rise to an energetically separated spectrum. We thus set the atom-laser detuning to be uniform across the lattice, $\delta_{i}=\delta$, and resonant with the target state, while keeping the driving Rabi frequency weak with respect to the energy gap with off-resonant states to minimize their unwanted excitation. We then compute the average nearest-neighbor concurrence and the state fidelity, $F=\langle\phi_{\rm RVB}|\rho_{st}|\phi_{\rm RVB}\rangle$, for the steady state $\rho_{st}$ predicted by the master equation (1). This analysis is performed for both the 2D waveguide and free-space scenarios while varying the atom-laser detuning. The results are plotted in Fig. 6 for different lattice configurations.
@@
These results demonstrate that certain features of the quantum dimer covering state can be probed in the steady state for small system sizes via a simple pumping scheme. As the system size increases, we expect that exciting the target state will become progressively more challenging. This difficulty arises both from the decreasing of the spacing between the energy levels in the spectrum, which reduces the spectral isolation of the target state, and from the fact that this state acquires a finite decay rate, causing it to rapidly decay once excited. Whether dimerization persists in the bulk of larger lattices thus remains an open question. In particular, in the 2D waveguide scenario, dissipation is expected to occur primarily through the edges of the lattice. Thus, even if the RVB state is not retained across the entire lattice, it could still dominate the steady state in the bulk region.

### v4

@@ II.3 State characterization
Figure 4: (a) Absolute value of the spin-spin correlation function for two lattices of sizes $N=2\times 10$ and $N=5\times 4$. In the first geometry we fixed: $\mathbf{i}=(1,1)$ and $\bm{\ell}=(1,\ell)$ while in the second $\mathbf{i}=(2,1)$ and $\bm{\ell}=(2,\ell)$. (b) Bipartite entanglement entropy in a ladder as a function of system size for two different partitions as indicated in the cartoon sketches. (c) Average nearest-neighbor concurrence as a function of inter-atomic distance for a $N=4\times 4$ lattice. In all plots, we considered the two models given in Eqs. (9) and (10) fixing the inter-atomic distance to $k_{0}d=0.1\pi$ and $k_{0}d=0.42\pi$, respectively. The results obtained from the RVB ansatz are reported with dashed line for direct comparison.
@@
    where $\lambda_{i}$ are the eigenvalues in decreasing order of the operator $R=\sqrt{\sqrt{\varrho}\tilde{\varrho}\sqrt{\varrho}}$ with $\tilde{\varrho}=(\sigma_{y}\otimes\sigma_{y})\varrho^{*}(\sigma_{y}\otimes\sigma_{y})$ and $\varrho$ is the corresponding density matrix. This measurement ranges from 0 for non-entangled states to 111 for maximally entangled configurations. Since an RVB liquid ansatz consists of a covering of nearest-neighbor dimers, we expect the two-qubit concurrence for this state to be finite only between nearest-neighbor sites. To quantify this, we define the average nearest-neighbour concurrence
@@
    where $N_{\langle ij\rangle}$ is the number of nearest neighbour atoms and $C(\varrho_{\langle ij\rangle})$ is the two-qubit concurrence computed for the density operator reduced to a given nearest neighbour pair $\langle ij\rangle$. The dependence of this quantity from the inter-atomic distance is shown in Fig. 4(c). The figure shows that there exists certain ranges of interatomic distances, consistent with those highlighted in Fig. 2, where the average nearest-neighbor concurrence remains close to the value predicted by the RVB ansatz (dashed line). This behavior confirms the formation of a dimer-like entanglement structure within the system for specific interatomic distance ranges.
@@ III Steady state preparation
Figure 5: (a) Scaling with system size of the decay rate and (b) of the spectral gap, $\Delta E$, of the least radiant state, $\gamma_{s=1}$, for model (9) (first panel) and model (10) (second panel). In the first case, the interatomic distance is fixed at $k_{0}d=0.1\pi$, while in the second case, it is set to $k_{0}d=0.42\pi$. In all cases, we compare two different lattice geometries and in (a) we included the decay rate scaling associated with $N/2$ dimers.
In this section, we explore possible strategies to probe the RVB state discussed in the previous section. In Sec. III.2, we discuss the finite lifetime and spectral isolation of the target state, which hinder its preparation via relaxation dynamics. In Sec. III, we instead propose a steady-state scheme that enables probing of the RVB state for small system sizes.
III.1 Lifetime and spectral isolation of the dimerized state
@@
As previously discussed, the least radiant state at half-filling is an eigenstate of the non-Hermitian Hamiltonian given in Eq. (8) and thus it has a finite decay rate. The scaling of this decay rate with the system size for the two cases in Eqs. (9)-(10) is shown in Fig. 5(a) for two different geometrical configurations. In all cases, we observe that the decay rate is upper-bounded by the single dimer decay rate times the number of excited emitters, $\gamma_{D}N/2$, where $\gamma_{D}$ is the decay rate of a single dimer, and remains relatively close to this bound. This implies that as the system size increases, even though this state remains the least radiant within the half-filling subspace, its decay rate can surpass that of an independent single emitter. This presents a significant challenge for preparing this state via a relaxation process. An initial half-filled configuration will rapidly radiate into excitation subspaces with lower filling, and the desired state will only be prepared with high probability under the condition that $N/2$ excitations remain throughout the system’s evolution, an event that becomes exponentially unlikely in time.
Another potential limitation for the preparation of the target RVB state arises from its spectral isolation. To quantify this, Fig. 5(b) shows the energy difference between the least radiant state in the half-filling sector and the closest state in energy within the same sector, plotted as a function of system size. In the 2D waveguide scenario, we observe a gradual decrease of this spectral gap with increasing system size, whereas in the free-space case, the gap remains comparatively stable. Based on these observations, we expect that the primary limitation in preparing the target state is related to its finite lifetime. This is confirmed in the following section, where we propose a possible driving scheme to address this challenge.
III.2 Steady-state protocol
Another possible approach to prepare the dimerized state is to employ a steady state protocol. In Ref. [58], two minimal conditions were identified for realizing a dimerized steady state: (i) the state must be completely dark, i.e., it must have a zero imaginary part in its eigenvalue, or equivalently, be annihilated by all jump operators, and (ii) it must be an eigenstate of the full coherent Hamiltonian, which includes both the photon-mediated interactions and the driving term. In the one-dimensional waveguide case, these conditions can be met under specific driving patterns, leading to a fully dark steady state composed of nearest-neighbor dimers (as given in Eq. (13)). However, in higher dimensions, these two conditions generally cannot be fulfilled simultaneously. In particular, in our case, as discussed above, the target state is not dark to begin with.
@@ III.2 Steady-state protocol
We thus employ a different steady-state protocol that targets the dimerized state by matching the energy and symmetry of the many-body wavefunction with the frequency and symmetry pattern of a driving field [95, 96, 97, 98]. More specifically, we assume that each atom in the array can be individually driven by a coherent field. This is incorporated by adding the following driving Hamiltonian (in the rotating wave approximation) to the master equation given in Eq. (7):
@@
where $\Omega_{i}$ is the driving strength on the $i$-th atom and $\delta_{i}$ is the detuning between the laser and the $i$-th atomic frequency. The pattern of the driving strength is chosen to have staggered (opposite) phases over the lattice to match the symmetry of the target many-body wave-function. This choice is motivated by the fact that, driving two atoms with opposite phase enhances the overlap with the singlet state, as shown in Ref. [96] for a 1D waveguide. Since the RVB target state is a superposition of singlet configurations, one might expect that applying a staggered driving phase across the array on a larger ensembles of emitters would yield a finite overlap with the desired state. This expectation is indeed confirmed by our simulations a posteriori, as shown in the following. For a finite-sized array, dipole-dipole interactions give rise to an energetically separated spectrum. We thus set the atom-laser detuning to be uniform across the lattice, $\delta_{i}=\delta$, and resonant with the target state, while keeping the driving Rabi frequency weak with respect to the energy gap with off-resonant states to minimize their unwanted excitation. We then compute the average nearest-neighbor concurrence and the state fidelity, $F=\langle\phi_{\rm RVB}|\rho_{st}|\phi_{\rm RVB}\rangle$, for the steady state $\rho_{st}$ predicted by the master equation (7). This analysis is performed for both the 2D waveguide and free-space scenarios while varying the atom-laser detuning. The results are plotted in Fig. 6 for different lattice configurations.
@@
These results demonstrate that certain features of the quantum dimer covering state can be probed in the steady state for small system sizes via a simple pumping scheme. As the system size increases, we expect that exciting the target state will become progressively more challenging.
Given the scaling of the spectral isolation of the target state shown in Fig. 5(b), this difficulty appears to be mainly attributed to the fact that the state acquires a finite decay rate, causing it to decay rapidly once excited. This also explains why state preparation in the free-space scenario performs significantly worse: the target state’s lifetime in that case is approximately an order of magnitude shorter than in the 2D waveguide setup.
Finally, we note that the relative robustness of the concurrence with increasing system size suggests that dimerization likely persists in the steady state, not as a coherent superposition of all dimer configurations, but rather as a mixed state composed of multiple dimer arrangements. In particular, we could expect such dimerization to better persists in the bulk of larger lattices. Indeed, in the 2D waveguide scenario, dissipation is expected to occur primarily through the edges of the lattice. Thus, even if the RVB state is not retained across the entire lattice, it could still dominate the steady state in the bulk region.